OpenStudy (sparklestaraa):

David's company reimburses his expenses on food, lodging, and conveyance during business trips. The company pays $60 a day for food and lodging and $0.65 for each mile traveled. David drove 600 miles and was reimbursed $3,390. Part A: Create an equation that will determine the number of days x on the trip Part B: Solve this equation justifying each step with an algebraic property of equality. Part C: How many days did David spend on this trip?

OpenStudy (sparklestaraa):

@dtan5457

OpenStudy (dtan5457):

First you need to make an equation. let x represent the amount of days 60x+ $0.65 for each mile traveled, 600 miles traveled 600 times 0.65=390 60x+390=3390

OpenStudy (dtan5457):

that is for part a

OpenStudy (dtan5457):

for b and c just solve x<<the amount of days

OpenStudy (dtan5457):

questions?

OpenStudy (sparklestaraa):

hold on I'm trying to figure out b & c xd

OpenStudy (sparklestaraa):

whats after 60 x + ?? and is the final equation 60x+390=3390

OpenStudy (dtan5457):

yes ^

OpenStudy (sparklestaraa):

so to find x i subtract 3390-390= 3000 60x=3000 then 3000/60 and x=50??

OpenStudy (dtan5457):

yes :)

OpenStudy (sparklestaraa):

and how do i justify it with an algebraic property of equality

OpenStudy (dtan5457):

60x+390=3390 when you subtract your 390 to both sides. it is the subtraction property of equality

OpenStudy (dtan5457):

like if 5=5 you subtract 2 on both sides. it will both equal 3.. subtraction property equality

OpenStudy (sparklestaraa):

how do i subtract from both sides? thats what I'm confused on a lot in math is the subtracting from both sides because don't i just need to subtract from one side?

OpenStudy (dtan5457):

you did it yourself lol look 60x+390=3390 you minus 390 on both sides theoretically 60x+390=3390 -390 -390 60x=3000

OpenStudy (dtan5457):

moar questions? <:)